# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v2

@@ Abstract.
Recent advancements in general-purpose AI have highlighted the importance of guiding AI systems towards the intended goals, ethical principles, and values of individuals and groups, a concept broadly recognized as alignment. However, the lack of clarified definitions and scopes of human-AI alignment poses a significant obstacle, hampering collaborative efforts across research domains to achieve this alignment. In particular, ML- and philosophy-oriented alignment research often views AI alignment as a static, unidirectional process (i.e., aiming to ensure that AI systems’ objectives match humans) rather than an ongoing, mutual alignment problem (Wikipedia, 2024). This perspective largely neglects the long-term interaction and dynamic changes of alignment. To understand these gaps, we introduce a systematic review of over 400 papers published between 2019 and January 2024, spanning multiple domains such as Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML), and others. We characterize, define and scope human-AI alignment. From this, we present a conceptual framework of “Bidirectional Human-AI Alignment” to organize the literature from a human-centered perspective. This framework encompasses both 1) conventional studies of aligning AI to humans that ensures AI produces the intended outcomes determined by humans, and 2) a proposed concept of aligning humans to AI, which aims to help individuals and society adjust to AI advancements both cognitively and behaviorally. Additionally, we articulate the key findings derived from literature analysis, including discussions about human values, interaction techniques, and evaluations. To pave the way for future studies, we envision three key challenges for future directions and propose examples of potential future solutions.
@@ 1. Introduction
Artificial Intelligence (AI) has advanced significantly, especially with the advent of general-purpose generative AI, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more (Rae et al., 2021; Chowdhery et al., 2023; Ouyang et al., 2022). However, as AI becomes increasingly powerful and deeply integrated into our lives, it also presents potential risks to people and to society broadly (Morris et al., 2024). For example, text-to-image generative models were found to amplify stereotypes about race and gender (blo, 2023), and biased algorithms in hiring processes were found to perpetuate discrimination (Park et al., 2021a). These risks highlight foundational questions regarding how and which values are embedded to create AI models that drive decisions within real-world contexts.
@@
Achieving this ongoing and mutual process of human-AI alignment necessitates a holistic understanding of the technical capabilities and limitations of AI, human-AI interaction, cognitive and social science, psychology and ethics, cross-cultural studies, and many other areas (Carroll et al., 2024). This paper contributes to the literature on alignment by presenting a comprehensive framework of this ongoing and mutual human-AI alignment process, creating a shared vocabulary among interdisciplinary communities, and envisioning future directions to achieve this long-term and dynamic alignment goal. To this end, we conduct a systematic literature review of over 400 papers across multiple disciplines, based on the PRISMA guidelines (Page et al., 2021; Stefanidi et al., 2023), present clarified definitions and scope (Section 3), derive a “Bidirectional Human-AI Alignment” framework shown in Figure 1 (Section 4) grounded to iterative paper coding and analysis (Section 5), and provide visions and potential solutions on three challenges for future research directions (Section 6).
Specifically, our systematic review takes an interdisciplinary perspective on human-AI alignment (Gilbert, 2024), drawing from theories across domains and empirical studies spanning model development and evaluation to interactive system design to human understanding and critical thinking to assess AI’s impact on individuals and society. The papers we reviewed in this survey are published around the advent of general-purpose generative AI, i.e., primarily between January 2019 and January 2024. These papers are selected to represent a holistic view of the human-AI alignment space, specifically from the domains of Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML), and others. Also, we form an interdisciplinary team to conduct the systematic review, consisting of researchers from the domains of HCI, NLP, ML, Data Science, Computational Social Science, and Cognitive Science.
@@
Consequently, we provide clarified definitions and scopes related to human-AI alignment, including “what is the goal of alignment?”, “with whom to align?”, and “what values should be aligned with?” (Section 3). We then present a conceptual framework of Bidirectional Human-AI Alignment from a long-term and dynamic perspective, encompassing both “Align AI to Humans” and “Align Humans to AI” (see Figure 1). The “Align AI to Humans” direction represents the conventional unidirectional studies that aim to integrate human specifications to train, steer, and customize AI. Crucially, our bi-directional framework places equal emphasis on the significance of “Align Humans to AI”, which represents humans’ cognitive and behavioral adaptation to the AI systems. The studies in this direction aim to support individuals and the broader society in understanding, critiquing, collaborating with, and adapting to transformative AI advancements. Furthermore, we identify four key research questions (RQs) in the Bidirectional Human-AI Alignment framework and provide a structured way to organize the existing research literature to address these questions in Section 4. The resulting structured topologies in Figure 6 and 7 aim to provide a shared vocabulary that can help streamline communication and collaboration between alignment researchers in different disciplines. Furthermore, through iterative paper coding and literature analysis, we derive insights including findings of human values, the interaction techniques for human-AI alignment, and the critical differences between human and AI evaluation (Section 5). To pave the way for future studies, we further envision three challenges in Figure 10, from near-term to long-term perspective, to motivate a dynamically co-evolving human-AI alignment. For each challenge, we articulate the future research problems and examples of potential future solutions from both directions in Section 6. Overall, we summarize our main contributions as follows:
Providing clarified definitions and scopes of human-AI alignment (i.e., including with whom to align, what is the alignment goal, and what are the values to be aligned with), and systematically reviewing over 400 relevant research papers (Section 3).
Developing the “Bidirectional Human-AI Alignment” framework to represent the ongoing, mutual alignment process and providing fine-grained taxonomies to address the involved four critical research questions (Section 4).
Discussing the key findings derived from literature analysis, including the human values and interaction techniques for alignment, and the discrepancy between AI and human evaluation (Section 5).
2. Background: A Historical View
@@
Outer and Inner Alignment. In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences (Wooldridge, 1999; Stuurman and Wijnands, 2001; Russell and Norvig, 2016). At the near-term stage, aligning AI involves two main challenges: carefully specifying111A specification often refers to a set of documented requirements, and/or particular information within them, to be satisfied by a material, design, product, or service. A specification is often a type of technical standard (International, 2012). the purpose of the system (outer alignment i.e., providing well-specified rewards (Ngo et al., 2024)) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals (Ngo et al., 2024)). Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) (Ouyang et al., 2022; Kirchner et al., 2022; Bai et al., 2022b, a) and to evaluate the performance of alignment (Santurkar et al., 2023). However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
@@
Specification Gaming. To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended (Hemphill, 2020). Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers (Wikipedia, 2024), which results in “specification gaming” (Krakovna et al., 2020) or “reward hacking” (Pan et al., 2022) issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways). One typical example of specification gaming shows that a simulated robot was trained to grab a ball by rewarding the robot for getting positive feedback from humans, but it learned to place its hand between the ball and camera, making it falsely appear successful (Christiano et al., 2017).222See the video at: https://en.wikipedia.org/wiki/File:Robot_hand_trained_with_human_feedback_%27pretends%27_to_grasp_ball.ogg Another classic (though extreme) example of an AI alignment problem is illustrated in the “paperclip maximizer” thought experiment: the scenario describes an advanced artificial intelligence tasked with manufacturing paperclips. If such a machine were not programmed to value human life, given enough power over its environment, it would try to turn all matter in the universe, including human beings, into paperclips or machines that manufacture further paperclips (Bostrom, 2020). Additionally, the black-box nature of neural networks further brings more ethical and safety concerns for alignment because humans cannot interpret the inner states and actions AI leveraged to achieve the final goals. Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern. Society is already facing these issues, such as data privacy (Tucker et al., 2018), algorithmic bias (Hajian et al., 2016), self-driving car accidents (Bonnefon et al., 2016), and more. As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective (Weidinger et al., 2023).
@@
Scalable Oversight. From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI (Morris et al., 2024)), it becomes increasingly difficult to align them through human feedback. Evaluating complex AI behaviors in increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with humans (Terry et al., 2023). Therefore, researchers have begun to investigate how to reduce the time and effort needed for human supervision, and how to assist human supervisors, referred to as Scalable Oversight (Amodei et al., 2016).
Existential Risk. Further, some AI researchers claim that (Bengio et al., 2024) advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training. For example, some language models seek power in text-based social environments by gaining money, resources, or social influence (Pan et al., 2023). Russell (2022) imagined a robot tasked to fetch coffee and evade shutdown since “you can’t fetch the coffee if you’re dead.” Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans (Dafoe and Russell, 2016).
@@
Dynamic Nature. As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well. As Dautenhahn et al. (2000) posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition). All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
3. Systematic Review Methodology
In this section, we aim to develop a holistic framework encompassing the ongoing and mutual process of human-AI alignment given the background in Section 2. To this end, we first introduce the core definitions of human-AI alignment (Section 3.1) and present the scope and its generalizability (Section 3.2). Next, we perform a systematic literature review guided by the PRISMA workflow (Section 3.3) and undertake an iterative coding process to achieve our bidirectional human-AI alignment framework (Section 4).
3.1. Definitions
To achieve the aforementioned ongoing and mutual process of human-AI alignment, we clarify three core questions before presenting the scope and process of systematic literature review: “With whom to align?”, “What is the goal of alignment?”, and “What values should be aligned with?”
3.1.1. With whom to align?
Pertinent stakeholders within the AI landscape can be the potential objects for AI to align with, including lay people (e.g., end users of AI systems) (Qi et al., 2023), AI practitioners (e.g., developers, researchers) (Boggust et al., 2022; Anwar et al., 2024; Ngo et al., 2024), organizational entities (e.g., technology firms, professional communities), national/international bodies (e.g., governments, legislative bodies) (Deshpande and Sharp, 2022) and others. Many alignment research papers have focused on general humans without specifying particular groups (Yao et al., 2023d; Liu et al., 2022b; Guo et al., 2024). Nevertheless, different groups hold different, sometimes even contrasting, values (Gabriel, 2020). As a consequence, rather than identifying a true moral theory as a one-size-fits-all value, prior alignment research argues to select the appropriate principles for compatible human groups (Sorensen et al., 2024). To this end, pluralistic value alignment, grounded on social choice theory (Arrow, 2012), proposes combining individual views fairly in developing alignment principles as a potential solution (Sorensen et al., 2024). In this work, we also consider values from this pluralistic perspective, where AI should be aligned with pluralistic human individuals and societal groups who would ultimately be impacted by AI.
3.1.2. What is the goal of alignment?
@@
Note that humans might not expect AI to encode all these values (e.g., social power), which could potentially result in harm and risks.
3.1.3. What are the values to be aligned with?
@@
Nevertheless, this conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment. Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature (elaborated in Section 3.3), mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content.
As a result, we identified the structural relationships among human values (see Figure 3) and mapped existing literature to a fine-grained taxonomy (see Figure 4).
As shown in Figure 3, we supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”).
@@
Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Figure 4.
@@
3.2. Scope
We define the scope of our literature review by specifying the key components we considered in human-AI alignment. We focus on a representative scope of alignment studies and discuss the generalizability of the scope after that.

### v3

@@ Abstract.
Recent advancements in general-purpose AI have highlighted the importance of guiding AI systems towards the intended goals, ethical principles, and values of individuals and groups, a concept broadly recognized as alignment. However, the lack of clarified definitions and scopes of human-AI alignment poses a significant obstacle, hampering collaborative efforts across research domains to achieve this alignment. In particular, ML- and philosophy-oriented alignment research often views AI alignment as a static, unidirectional process (i.e., aiming to ensure that AI systems’ objectives match humans) rather than an ongoing, mutual alignment problem (Wikipedia, 2024). This perspective largely neglects the long-term interaction and dynamic changes of alignment. To understand these gaps, we introduce a systematic review of over 400 papers published between 2019 and January 2024, spanning multiple domains such as Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML). We characterize, define and scope human-AI alignment. From this, we present a conceptual framework of “Bidirectional Human-AI Alignment” to organize the literature from a human-centered perspective. This framework encompasses both 1) conventional studies of aligning AI to humans that ensures AI produces the intended outcomes determined by humans, and 2) a proposed concept of aligning humans to AI, which aims to help individuals and society adjust to AI advancements both cognitively and behaviorally. Additionally, we articulate the key findings derived from literature analysis, including literature gaps and trends, human values, and interaction techniques. To pave the way for future studies, we envision three key challenges and give recommendations for future research.
@@ 1. Introduction
Artificial Intelligence (AI) has advanced significantly, especially with the advent of general-purpose generative AI, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more (Ouyang et al., 2022). However, as AI becomes increasingly powerful and integrated into daily life, society is confronted with an array of risks to both individuals and to broader society (Morris et al., 2024). For example, text-to-image generative models were found to amplify stereotypes about race and gender (blo, 2023), and biased algorithms in hiring processes were found to perpetuate discrimination (Park et al., 2021a). These risks illuminate foundational questions around how and which values are embedded in AI models that drive decisions within real-world contexts.
@@
This paper conducts a systematic literature review of over 400 human-AI alignment papers across multiple disciplines, based on PRISMA guidelines (Page et al., 2021; Stefanidi et al., 2023). The literature review corpus includes papers drawn from the the Human-Computer Interaction (HCI), Natural Language Processing (NLP) and Machine Learning (ML) domains that were published between the advent of general-purpose generative AI to present, i.e., primarily between January 2019 and January 2024. We take an interdisciplinary approach to human-AI alignment (Gilbert, 2024), drawing from theories and empirical studies across the aforementioned domains. To round out our approach, we composed our research team of members from the domains of HCI, NLP, ML, Data Science, Computational Social Science, and Cognitive Science.
| Definition. Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI to Humans’ and ‘Aligning Humans to AI’. The former focuses on integrating human specifications to train, steer, and customize AI, while the latter investigates human cognitive and behavioral adaptations to AI, which supports humans in understanding, critiquing, collaborating with, and adapting to AI advancements. |
@@
Paper Organization and Key Findings. We provide clarified definitions and scopes related to human-AI alignment, including “what is the goal of alignment?”, “with whom to align?”, and “what values should be aligned with?”(Section 3). We then present a conceptual framework of Bidirectional Human-AI Alignment from a long-term and dynamic perspective, encompassing both “Align AI to Humans” and “Align Humans to AI” (see Figure 1). Furthermore, we identify four key research questions (RQs) in the Bidirectional Human-AI Alignment framework and provide a structured way to organize the existing research literature to address these questions in Section 4. The resulting structured topologies in Figure 2 and 3 aim to provide a shared vocabulary that can help streamline communication and collaboration between alignment researchers in different disciplines. Furthermore, through iterative paper coding and literature analysis, we derive insights including meta analysis on gaps and trends in existing alignment research, findings of human values, and the interaction techniques for human-AI alignment (Section 5). To pave the way for future studies, we further envision three future directions in Figure 6 and Section 6. Our study highlights several key findings. For instance, many essential human values outlined in Table 2, such as loyalty and environmental protection, are crucial for alignment but are often overlooked in current research. Additionally, many topics within the bidirectional alignment framework, including leveraging implicit human feedback to align AI models and auditing AI for various social values, are important but underexplored.
Operationalizing the Framework. To help AI practitioners (e.g., developers, researchers, user experience designers) who are interested in this field to leverage the bidirectional approaches, we highlight some systematic resources about human-AI alignment provided both within this survey and on companion Github repository:
A Github repository of “Bidirectional Human-AI Alignment” maintaining latest reading lists and relevant updates.
Fig. 1: An overview of the bidirectional human-AI alignment framework with formal definition.
Table 1: Typical alignment goals, their definitions, and limitations.
Table 2: A comprehensive list of basic human values, their relationships, and related alignment papers.
Fig. 4: An illustration of trends and gaps of alignment research dimension distribution.
Fig. 5: A summary of common interactive techniques to specify human values and exemplary papers.
2. Challenges in Achieving Alignment
@@
Challenge 1: Outer and Inner Alignment. In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences (Wooldridge, 1999; Stuurman and Wijnands, 2001; Russell and Norvig, 2016). At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards (Ngo et al., 2024)) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals (Ngo et al., 2024)). Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) (Ouyang et al., 2022) and to evaluate the performance of alignment (Santurkar et al., 2023). However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
@@
Challenge 2: Specification Gaming. To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended (Hemphill, 2020). Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers (Wikipedia, 2024), which results in “specification gaming” (Krakovna et al., 2020) or “reward hacking” (Pan et al., 2022) issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways). Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output. Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern. Society is already facing these issues, such as data privacy (Tucker et al., 2018), algorithmic bias (Hajian et al., 2016), self-driving car accidents (Bonnefon et al., 2016), and more. As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective (Weidinger et al., 2023).
@@
Challenge 3: Scalable Oversight. From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI (Morris et al., 2024)), it becomes increasingly difficult to align them to human values through human feedback. Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values (Terry et al., 2023). Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight (Amodei et al., 2016).
@@
Challenge 4: Dynamic Nature. As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well. As Dautenhahn et al. (2000) posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition). All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
Challenge 5: Existential Risk. Further, some AI researchers claim that (Bengio et al., 2024) advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training. For example, some language models seek power in text-based social environments by gaining money, resources, or social influence (Pan et al., 2023). Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans (Dafoe and Russell, 2016).
3. Fundamental Definitions and Clarifications
Addressing alignment challenges is a complex and multifaceted process. To develop a comprehensive perspective of the ongoing and mutual process of human-AI alignment, this section begins by introducing core definitions and key components of alignment. Subsequently, we outline our methodology for conducting a systematic literature review and the rationale behind our proposed framework of bidirectional human-AI alignment.
3.1. What is the goal of alignment?
@@
3.2. With whom to align?
Pertinent stakeholders within the AI landscape can be the potential objects for AI to align with, including lay people (e.g., end users of AI systems) (Qi et al., 2023), AI practitioners (e.g., developers, researchers) (Boggust et al., 2022; Anwar et al., 2024; Ngo et al., 2024), organizational entities (e.g., technology firms, professional communities), national/international bodies (e.g., governments, legislative bodies) (Deshpande and Sharp, 2022) and others. Many alignment research papers have focused on general humans without specifying particular groups (Yao et al., 2023b; Liu et al., 2022; Guo et al., 2024). Nevertheless, different groups hold different, sometimes even contrasting, values (Gabriel, 2020). As a consequence, rather than identifying a true moral theory as a one-size-fits-all value, prior alignment research argues to select the appropriate principles for compatible human groups (Sorensen et al., 2024). To this end, pluralistic value alignment, grounded on social choice theory (Arrow, 2012), proposes combining individual views fairly in developing alignment principles as a potential solution (Sorensen et al., 2024). In this work, we also consider values from this pluralistic perspective, where AI should be aligned with pluralistic human individuals and societal groups who would ultimately be impacted by AI.
3.3. What are the values to be aligned with?
@@
A Comprehensive Taxonomy of Human Values.
Nevertheless, this conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment. Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature (elaborated in Section A.1), mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content.
As a result, we identified the structural relationships among human values (see Figure LABEL:fig:value_structure) and mapped existing literature to a fine-grained taxonomy (see Table 2).
As shown in Figure LABEL:fig:value_structure, we supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”).
@@
Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 2.
@@
| Takeaways & Implications for Alignment Clarifications 1. We propose the “pluralistic human values” as the alignment goal, wherein AI should be aligned with the diverse perspectives of individuals and societal groups who may be directly or indirectly impacted by AI. 2. Many exemplary human values outlined in Table 2, such as loyalty and environmental protection, are important for alignment but are often overlooked or underexplored in current research. 3. The human value system should account for complex application contexts and relationships, such as priorities and conflicts, rather than focusing on an exclusive subset of values. |
3.4. Scopes and Key Components in Alignment
This section clarifies the key components we considered in human-AI alignment. Additionally, we define the scopes of each key component as commonly adopted in current research.